# SincVAE: a New Approach to Improve Anomaly Detection on EEG Data Using SincNet and Variational Autoencoder

### v1

@@ 3 Proposed Method
In this section, we will introduce the SincNet model to the reader, followed by an overview of the VAE framework. Finally, we will present SincVAE, the proposed method of this work.
@@ 3.1 SincNet
where is the input signal segment, is the filter of length , denotes the convolution operation, and is the output. Typically, each element of the filter is learned during the training phase.
SincNet modifies this process by using a pre-defined function that depends on a limited number of learnable parameters :
This function is designed to implement a filterbank consisting of rectangular bandpass filters. The magnitude of a generic bandpass filter, in the frequency domain, can be expressed as the difference between two low-pass filters as follows [50]:
@@
The cutoff frequencies are initialized randomly within the range , where is the sampling frequency of the input signal. To ensure that and , the parameters are adjusted as:
The authors in [50] points out that the training process naturally keeps below the Nyquist frequency [76], eliminating the need for explicit constraints.
Additionally, SincNet applies a windowing function [77] to smooth the discontinuities at the ends of . This is achieved by multiplying with a Hamming window [78]:
where is defined as:
@@ 3.2 Variational Autoencoder
A Variational Autoencoder (VAE) is a generative model made of two principal components: a probabilistic decoder and a probabilistic encoder. The decoder, with parameters , models the likelihood function and generates new data given a latent variable . The encoder, with parameters , models the posterior distribution to approximate the true intractable posterior [34]. During the training of a VAE, both and are optimized through the following generative process [79]:
which can be re-written as:
where is the non-negative Kullback–Leibler (KL) divergence and represents the prior distribution over the latent variables [79]. Since the KL divergence is non-negative, the term is called Evidence Lower Bound (ELBO) of , and it can be rewritten as below:
where the second term is an expected reconstruction error between the input and generated data.
The VAE optimization can be focused on maximizing the ELBO [79]. A key challenge in this process is the need to sample random latent variables from , which makes the training intractable. The reparametrization trick is employed to avoid this problem: by assuming both the prior and the posterior to be Gaussian distributions with a diagonal covariance matrix, with the prior set to the isotropic unit Gaussian , each random variable is reparametrized as differential transformation of a noise variable as follows [79]:
Under this framework, the ELBO can be differentiated and optimized with respect to the parameters and . Specifically, the ELBO can be maximized using gradient based methods. This approach allows for considerable flexibility in the design of both the probabilistic encoder and the probabilistic decoder.
@@ 3.3 SincVAE
This work proposes SincVAE, which integrates SincNet as the first layer in the VAE’s probabilistic encoder. This model is specifically designed to enhance anomaly detection in time series data. In particular, this design is aimed at learning an ad-hoc array of bandpass filters, facilitating the decomposition of the input time series and enhancing the feature extraction within the time series processing.
The reconstruction error, computed in terms of Mean Squared Error (MSE) [27], between the input data and its reconstruction is involved into the classification process given a threshold . A graphical representation of the proposed pipeline is shown in Figure 1.
@@
This work explores the impact of SincVAE on anomaly detection challenges within the healthcare sector, with a focus on seizure detection in EEG data. It is expected the integration of SincNet within the VAE framework in the seizure detection problem can enhance the learning of normal patterns within EEG data, thus leading to a more robust identification of anomalous patterns. The SincVAE’s ability to learn custom bandpass filters could not only to boost the overall efficacy of the seizure detection process but also to potentially reduce or even eliminate the preprocessing currently required to identify band frequencies where informative content is more present.
@@ 4 Experimental Assessment
In this work, the effectiveness of SincVAE for seizure detection was analyzed through experimental results obtained from two datasets, i.e. the Bonn dataset [80] and the CHB-MIT dataset [81]. These datasets have been widely utilized in numerous studies, including [82, 83, 84, 85, 86, 87, 88, 70, 89]. The seizure detection problem was tackled with a semi-supervised approach, thus models were trained exclusively on non-seizure data. The classification of data into seizure or non-seizure categories is determined by analyzing the reconstruction error generated by the model.
For a robust comparative analysis, the experiments employed a fixed base VAE architecture in two distinct configurations: (i) the VAE with a SincNet layer on top of the encoder network, referred as SincVAE, and (ii) the VAE without the SincNet layer, referred as VAE. Specifically, the AE-CDNN architecture detailed in [69] was adopted, with modifications applied to the specific needs of VAE operations, particularly in the latent space and training configurations. This methodological choice facilitates a clear examination of the SincNet layer’s contribution to the VAE’s performance. In accordance with [50], the SincNet layer was followed by a Layer Normalization and an activation function.
@@ 4.1 The Bonn Dataset
Dataset description.
The Bonn dataset [80], acquired by the Bonn University in Germany, consists in five distinct collections of EEG signals. Each collection includes 100 single-channel EEG segments, each lasting 23.6 seconds, derived from continuous multi-channel EEG recordings. These segments were selected after a visual inspection to eliminate artifacts. Each EEG segment was captured using a 128-channel amplifier system paired with a 12-bit analog to digital converter at a sampling rate of . The dataset is organized into five groups:
Sets A and B contain EEG recordings from healthy volunteers. In particular, Set A contains data acquired during eyes-open condition, and Set B data acquired during eyes-closed conditions;
@@
Data preprocessing.
The final 20 % of each training dataset is kept apart for testing the models on non-seizure data instances. Then, the data were segmented into one-second frames, resulting in 1840 samples each with a length of 173 data points. Subsequently, the dataset underwent filtering with a 40 Hz low-pass filter following [80]. Z-score normalization [92] was then applied. A summary of the Bonn dataset, including details on the preprocessing and data partitioning, is presented in Table 1.
@@
Model selection.
The model selection stage was performed through the search space defined in Table 2.
@@
The total number of configuration explored was 2295. 10-fold Cross Validation [61] was employed as the resampling method, with the mean MSE across the 10 folds used as the criterion for selecting the best architecture. In this setup, for each fold of the Cross Validation, 20 % of the training data was sampled and designated as the validation set. This analysis was done separately on the Set A and Set B. For the sake of clarity, the methodology in this paragraph is explained through the results obtained on the Set A only.
From the Cross Validation analysis, it was observed that the configuration with the lowest mean MSE across the folds showed a distribution of results that overlapped with several other configurations . This indicated that while this configuration performed well in terms of MSE, it was not distinctly superior to others. Thus, other analyses were done.
@@
Then, the normality of the results for each configuration was evaluated using the Shapiro-Wilk test [93]. By setting the significance level at , it was found that 42 out of the 244 configurations yielded a p-value lower than , indicating that their result distributions do not conform to a normal distribution.
Consequently, the Kruskal-Wallis test [94] was applied to compare all configurations, using the same significance level . This test produced a p-value lower than , confirming significant differences among the distributions of the configurations.
Then, a pairwise comparison of the remaining configurations was conducted using the Mann-Whitney U-test [95]. The results are shown in Figure 2, where the 244 configurations are ordered in ascending order of mean MSE over the folds.
@@
As stated above, the best configuration was chosen prioritizing the lowest model complexity among these 131 configurations. Thus, the latent space was first fixed to 32, the Identity activation function was chosen as activation function of the SincNet layer, and 16 filters were selected with a kernel size of 41.
@@ 4.2 The CHB-MIT Dataset
| Subject | Gender | Age | # Tracks | # Seizures Tracks |
| 1 | F | 11 | 42 | 7 |
| 2 | M | 11 | 36 | 3 |
| 3 | F | 14 | 38 | 7 |
| 4 | M | 22 | 42 | 3 |
| 5 | F | 7 | 39 | 5 |
| 6 | F | 1.5 | 18 | 7 |
| 7 | F | 14.5 | 19 | 3 |
| 8 | M | 3.5 | 20 | 5 |
| 9 | F | 10 | 19 | 3 |
| 10 | M | 3 | 25 | 7 |
| 11 | F | 12 | 35 | 3 |
| 12 | F | 2 | 24 | 13 |
| 13 | F | 3 | 33 | 8 |
| 14 | F | 9 | 26 | 7 |
| 15 | M | 16 | 40 | 14 |
| 16 | F | 7 | 19 | 7 |
| 17 | F | 12 | 21 | 3 |
| 18 | F | 18 | 36 | 6 |
| 19 | F | 19 | 30 | 3 |
| 20 | F | 6 | 29 | 6 |
| 21 | F | 13 | 33 | 4 |
| 22 | F | 9 | 31 | 3 |
| 23 | F | 6 | 9 | 3 |
| 24 | N/A | N/A | 22 | 12 |
All EEG signals in the CHB-MIT dataset were recorded at a sampling rate of with a 16-bit resolution. These recordings conform to the International 10-20 system for EEG electrode placement and naming conventions. The number of channels recorded varied among subjects, with a minimum of 23 channels used. In the cases of subjects 4 and 9, the dataset includes additional channels for ECG and vagal nerve stimulus. In this work, the experiments are based only the data acquired from the 23 channels used by all the subjects. Subjects’ recordings are organized into tracks, which are labeled according to whether they contain seizure activity. For tracks with seizures, the specific time intervals of the seizure occurrences are meticulously documented111For further details on the CHB-MIT dataset, see https://physionet.org/content/chbmit/1.0.0/.
@@
Data were segmented into frames of one second of length 256 with 23 channels. Then, the dataset was filtered using a bandpass filter of , following [22]. Following this, Z-score normalization was applied. Finally, for each subject, a random sample consisting of 1-second windows totaling 10 minutes was drawn from the non-seizure tracks to test the models’ ability to detect non-seizure cases. The remaining data were used to train the models. Details on the CHB-MIT dataset are provided in Table 4.
| Subject | Training set | Test set |
| 1 | 114230 | 1042 |
| 2 | 110969 | 691 |
| 3 | 103777 | 1002 |
| 4 | 107686 | 760 |
| 5 | 114578 | 1158 |
| 6 | 156677 | 669 |
| 7 | 181777 | 925 |
| 8 | 46187 | 1519 |
| 9 | 208250 | 726 |
| 10 | 114604 | 1047 |
| 11 | 107368 | 1406 |
| 12 | 60628 | 664 |
| 13 | 133796 | 815 |
| 14 | 64182 | 727 |
| 15 | 89375 | 1944 |
| 16 | 38989 | 638 |
| 17 | 35390 | 893 |
| 18 | 96573 | 917 |
| 19 | 85776 | 836 |
| 20 | 78707 | 741 |
| 21 | 96572 | 799 |
| 22 | 92985 | 804 |
| 23 | 68109 | 713 |
| 24 | 38989 | 929 |
Model selection.
@@
To manage the increased computational effort of the model selection stage, necessitated by the larger dataset size, the search space was limited based on insights gained from the Bonn dataset analysis. Specifically, only larger kernel sizes were chosen, and the Tanh activation function was excluded due to its observed negative impact on performance. In particular, the total number of configuration explored was 210.
Similar to the results obtained on the Bonn dataset, the configuration with the lowest mean MSE across the tracks showed notable overlap with several other configurations. Thus, only those configurations whose mean MSE falls within one standard deviation of the lowest mean MSE were retained for further analysis. After this filtering, the top 67 configurations were selected for further detailed analysis.
By applying the Shapiro-Wilk test to the results of each configuration, and using a significance level of , none of the configurations rejected the null hypothesis, suggesting that the data from all configurations can be considered normally distributed. Consequently, an ANOVA test [96] was conducted using the same significance level. This test also did not reject the null hypothesis, indicating that there were no significant differences among the configurations. This result implies that the performance across different configurations is statistically comparable under the conditions tested.

### v2

@@ 3 Proposed Method
This section introduces the SincNet model, followed by an overview of the VAE framework. Finally, it presents SincVAE, the proposed method developed in this work.
@@ 3.1 SincNet
where is the input signal segment, is the filter of size , denotes the convolution operation, and is the output. Typically, each element of the filter is learned during the training phase. SincNet modifies this process by using a pre-defined function that depends on a limited number of learnable parameters :
The function is designed to implement a filterbank consisting of rectangular bandpass filters. The magnitude of a generic bandpass filter, in the frequency domain, can be expressed as the difference between two low-pass filters as follows [51]:
@@
The cutoff frequencies are randomly initialized within the range , where is the sampling frequency of the input signal. To ensure that and , the parameters are adjusted as:
The authors in [51] point out that the training process naturally keeps below the Nyquist frequency [85], eliminating the need for explicit constraints.
Additionally, SincNet applies a windowing function [86] to smooth the discontinuities at the boundaries of . This is achieved by multiplying with a window function :
where the windowing function is the Hamming window [87], defined as
@@ 3.2 Variational Autoencoder
A variational autoencoder (VAE) [35] is a generative model comprising a probabilistic decoder and encoder. The decoder , with parameters , generates new data given a latent variable , while the encoder , with parameters , approximates the true posterior , which is generally intractable. A prior , typically the isotropic unit Gaussian , is placed over the latent variables.
To train a VAE, the parameters and are optimized by maximizing the evidence lower bound (ELBO) of the marginal log-likelihood :
where is the Kullback-Leibler divergence. The first term represents the expected reconstruction accuracy, and the second term regularizes the latent space by ensuring is close to the prior . To make the optimization tractable, the reparameterization trick is employed. Assuming is a Gaussian with diagonal covariance, each latent variable is reparameterized as a deterministic transformation of a noise variable [88]:
This reparameterization enables backpropagation through stochastic sampling, making the ELBO differentiable with respect to the model parameters, and allowing it to be optimized using gradient-based methods. Under this framework, the ELBO can be maximized, offering significant flexibility in the design of the encoder and decoder architectures.
@@ 3.3 SincVAE
VAEs represent a versatile framework where both the encoder and decoder can be implemented in various ways. This flexibility allows the method to be adapted to different types of data. In particular, several research works have employed VAEs implemented with convolutional layers to process time series data (see, for example [89, 90]). During the training stage, filters of the convolutional layers are optimized to extract the features necessary to solve a given task. While these filters may be meaningful to the neural network, they are not intuitive for humans and, on time series, they can often take noisy and incongruous multi-band shapes [51]. Also, a crucial component of current waveform-based CNNs is the first convolutional layer, since it handles high-dimensional inputs and is particularly susceptible to vanishing gradient issues, especially in very deep architectures [51], and this aspect might affect the performance of the whole network. Thus, to simplify the feature extraction stage of the neural network, one approach could involve preprocessing the time series to reduce the complexity of the input data using various techniques, including signal decomposition via fast Fourier transform (FFT) [91], which is commonly applied in the BCI domain (see, for example [92, 93]). However, this approach is restricted to predefined frequency bands well-established in the literature or requires a prior analysis phase where experts examine the signal to identify points of interest. This process can be time-consuming and prone to errors, potentially affecting the overall accuracy in solving a given task. As stated by the authors in [51], SincNet has the potential to address these issues. Acting as band-pass filters with a minimal number of trainable parameters, SincNet filters in the first layer decompose the signal into frequency bands during the optimization process. This approach provides filters that can be interpreted by humans, eliminates the need for manual preprocessing to simplify the input data and enhances the feature extraction performed by subsequent convolutional layers.
In the context of anomaly detection using VAEs in a semi-supervised setting, the model is trained exclusively on non-anomalous data. The goal is to optimize the feature extraction and reconstruction processes - handled by the encoder and decoder networks, respectively - specifically for normal data instances [39]. As a result, once training is complete, the VAE’s encoder is expected to encounter difficulties in extracting meaningful features from anomalous data, leading to poor data reconstructions. In this framework, the more effectively the VAE performs feature extraction on normal data, the clearer the distinction between anomalous and non-anomalous inputs will be.
The proposed method, referred to as SincVAE, incorporates SincNet as a convolutional layer preceding the VAE’s probabilistic encoder. Specifically designed to improve anomaly detection on time series data, SincVAE leverages SincNet’s ability to learn a customized array of bandpass filters while providing a set of human-interpretable filters. This allows the decomposition of input time series, thereby enhancing the VAE’s feature extraction and improving the overall effectiveness of time series processing. In particular, the introduction of the SincNet layer is expected to improve the VAE’s functioning in extracting features more representative of a non-anomalous state during the training stage. Consequently, this should result in producing poorer reconstructions for anomalous data, thereby emphasizing the distinction between non-anomalous and anomalous data, and improving classification accuracy. Within this framework, the reconstruction error, computed in terms of MSE between the input data and its reconstruction, is involved in the classification process given a threshold . Specifically, an MSE whose value is higher than , will indicate that the input data is anomalous; conversely, it will indicate that the input data is non anomalous. A graphical representation of the proposed pipeline is shown in Figure 1.
@@
This study investigates the effectiveness of SincVAE in addressing anomaly detection challenges in the healthcare domain, with a specific focus on seizure detection from EEG data. It is expected that the integration of SincNet within the VAE framework in the seizure detection problem can enhance the learning of non-seizure patterns within EEG data, thus leading to a more robust identification of seizure patterns. The SincVAE’s ability to learn custom bandpass filters could not only boost the overall efficacy of the seizure detection process, but also potentially reduce (or, even, eliminate) the preprocessing often required to identify band frequencies where informative content is more present, which is frequently done in the BCI context.
@@ 4 Experimental Assessment
In this work, the effectiveness of SincVAE for seizure detection was analyzed through experimental results obtained from two datasets: the Bonn dataset [94] and the CHB-MIT dataset [95]. These datasets have been widely utilized in numerous studies, including [96, 97, 98, 99, 100, 101, 102, 79, 103]. The seizure detection problem was tackled with a semi-supervised approach, therefore, models were trained exclusively on non-seizure data. The classification of data into seizure or non-seizure categories is determined by analyzing the reconstruction error obtained from the input data and their reconstructions generated by the model.
For a robust comparative analysis, the experiments employed a fixed base VAE architecture in two distinct configurations: (i) the VAE with a SincNet preceding the encoder network, referred to as SincVAE, and (ii) the VAE without the SincNet layer, referred to as VAE. Specifically, the AE-CDNN architecture detailed in [78] was adopted, with modifications applied to the specific needs of VAE operations, particularly in the latent space and training configurations. This methodological choice facilitates a clear examination of the SincNet layer’s contribution to the VAE’s performance. In accordance with [51], the SincNet layer was followed by a Layer Normalization and an activation function.
@@ 4.1 The Bonn Dataset
Dataset description
The Bonn dataset [94], acquired by the Bonn University in Germany, consists of five distinct collections of EEG signals. Each collection includes 100 single-channel EEG segments, each lasting 23.6 seconds, derived from continuous multi-channel EEG recordings. These segments were selected after a visual inspection to eliminate artifacts. Each EEG segment was captured using a 128-channel amplifier system paired with a 12-bit analog to digital converter at a sampling rate of . The dataset is organized into five groups:
Sets A and B contain EEG recordings from healthy volunteers. In particular, Set A contains data acquired during eyes-opened condition, and Set B data acquired during eyes-closed conditions;
@@
Data preprocessing
First, the data were segmented into one-second frames, yielding 2300 samples of 173 data points each for Sets A, B, and E. In both cases (i.e., Set A vs Set E and Set B vs Set E), the final of each training dataset (Sets A and B) was set aside to evaluate the models’ ability to classify non-seizure data alongside the seizure data in Set E. This resulted in 1840 training samples and 2760 testing samples. Subsequently, the dataset underwent filtering with a 40 Hz low-pass filter following [94]. Z-score normalization [106] was then applied. A summary of the Bonn dataset, including details on the preprocessing and data partitioning, is presented in Table 1.
@@
Model selection
The model selection stage was performed through the search space defined in Table 2. As previously mentioned, the hyperparameter ranges were selected based on preliminary experiments. These initial trials allowed us to identify areas of the search space where training generally led to promising outcomes, which were subsequently used to define the grid for model selection.
@@
The total number of configuration explored was 2295. 10-fold cross validation [70] was employed as the resampling method, with the mean MSE across the 10 folds used as the criterion for selecting the best architecture. In this setup, for each fold of the cross validation, of the training data was sampled and designated as the validation set. This analysis was done separately on the Set A and Set B. For the sake of clarity, the methodology in this paragraph is explained through the results obtained on the Set A only.
From the cross validation analysis, as shown in Figure 2, the configuration with the lowest mean MSE across the folds exhibited a result distribution that overlapped with several other configurations. This suggests that, while this configuration demonstrated strong performance in terms of MSE, it was not distinctly superior, highlighting the need for further analysis through statistical hypothesis testing. To systematically assess performance differences among configurations, we followed a structured statistical testing procedure. First, the Shapiro-Wilk [107] test was applied to evaluate the normality of each configuration’s result distribution. This informed the choice between parametric (ANOVA [110]) and non-parametric (Kruskal-Wallis [108], Mann-Whitney U-test [109]) tests for subsequent comparisons. When statistical tests revealed significant differences, we prioritized configurations with superior performance as indicated by the test outcomes. Conversely, when no significant differences were found, configurations were considered statistically equivalent, and the final choice was guided by practical considerations such as model complexity or computational cost.
@@
Then, the normality of the results for each configuration was evaluated using the Shapiro-Wilk test. By setting the significance level at , it was found that 42 out of the 244 configurations yielded a p-value lower than , indicating that their result distributions do not conform to a normal distribution. Consequently, the Kruskal-Wallis test was applied to compare all configurations, using the same significance level . This test produced a p-value lower than , confirming significant differences among the distributions of the configurations.
Then, a pairwise comparison of the remaining configurations was conducted using the Mann-Whitney U-test. The results are shown in Figure 3, where the 244 configurations are ordered in ascending order of mean MSE over the folds.
@@
As mentioned earlier, the best configuration was selected by prioritizing the lowest model complexity among the 131 configurations. From these, the configurations with the smallest latent space, specifically 32, were chosen. Then, from the remaining options, the configuration with the Identity function as the activation function for the SincNet layer was selected. Finally, among the remaining configurations, those with 16 filters were chosen, leading to the final selection of a kernel size of 41.
@@ 4.2 The CHB-MIT Dataset
| Subject | Gender | Age | # Tracks | # Seizures Tracks | Training set | Test set |
| 1 | F | 11 | 42 | 7 | 114230 | 1042 |
| 2 | M | 11 | 36 | 3 | 110969 | 691 |
| 3 | F | 14 | 38 | 7 | 103777 | 1002 |
| 4 | M | 22 | 42 | 3 | 107686 | 760 |
| 5 | F | 7 | 39 | 5 | 114578 | 1158 |
| 6 | F | 1.5 | 18 | 7 | 156677 | 669 |
| 7 | F | 14.5 | 19 | 3 | 181777 | 925 |
| 8 | M | 3.5 | 20 | 5 | 46187 | 1519 |
| 9 | F | 10 | 19 | 3 | 208250 | 726 |
| 10 | M | 3 | 25 | 7 | 114604 | 1047 |
| 11 | F | 12 | 35 | 3 | 107368 | 1406 |
| 12 | F | 2 | 24 | 13 | 60628 | 664 |
| 13 | F | 3 | 33 | 8 | 133796 | 815 |
| 14 | F | 9 | 26 | 7 | 64182 | 727 |
| 15 | M | 16 | 40 | 14 | 89375 | 1944 |
| 16 | F | 7 | 19 | 7 | 38989 | 638 |
| 17 | F | 12 | 21 | 3 | 35390 | 893 |
| 18 | F | 18 | 36 | 6 | 96573 | 917 |
| 19 | F | 19 | 30 | 3 | 85776 | 836 |
| 20 | F | 6 | 29 | 6 | 78707 | 741 |
| 21 | F | 13 | 33 | 4 | 96572 | 799 |
| 22 | F | 9 | 31 | 3 | 92985 | 804 |
| 23 | F | 6 | 9 | 3 | 68109 | 713 |
| 24 | N/A | N/A | 22 | 12 | 38989 | 929 |
All EEG signals in the CHB-MIT dataset were recorded at a sampling rate of with a 16-bit resolution. These recordings conform to the International 10-20 system for EEG electrode placement and naming conventions. The number of channels recorded varied among subjects, with a minimum of 23 channels used. In the cases of subjects 4 and 9, the dataset includes additional channels for ECG and vagal nerve stimulus. In this work, the experiments were based only the data acquired from the 23 channels used by all the subjects. Subjects’ recordings were organized into tracks, which are labeled according to whether they contain seizure activity or not. For tracks with seizures, the specific time intervals of the seizure occurrences were meticulously documented by the dataset’s authors111For further details on the CHB-MIT dataset, see https://physionet.org/content/chbmit/1.0.0/.
@@
Data were segmented into frames of one second of length 256 with 23 channels. Then, the dataset was filtered using a bandpass filter of , following [23]. Following this, Z-score normalization was applied. Finally, for each subject, a random sample consisting of 1-second windows totaling 10 minutes was drawn from each non-seizure track to test the models’ ability to detect non-seizure cases. The remaining data were used to train the models. Details on the CHB-MIT dataset are provided in Table 3.
Model selection
@@
To manage the increased computational effort of the model selection stage, necessitated by the larger dataset size, the search space was limited based on insights gained from the Bonn dataset analysis. In particular, the total number of configuration explored was 210.
Similar to the results obtained on the Bonn dataset, the configuration with the lowest mean MSE across the tracks showed notable overlap with several other configurations, requiring additional analyses using statistical hypothesis testing. Thus, only those configurations whose mean MSE falls within one standard deviation of the lowest mean MSE were retained for further analysis. After this filtering, the top 67 configurations were selected for further detailed analysis.
By applying the Shapiro-Wilk test to the results of each configuration, and using a significance level of , none of the configurations rejected the null hypothesis, suggesting that the data from all configurations can be considered normally distributed. Consequently, an ANOVA test was conducted using the same significance level. This test also did not reject the null hypothesis, indicating that there were no significant differences among the configurations. This result implies that the performance across different configurations is statistically comparable under the conditions tested.
@@
4.3 Sensitivity analysis of SincVAE’s hyperparameters
This subsection examines the sensitivity of hyperparameters for the SincVAE architecture, building on the results from the model selection stages, which aimed to identify the optimal hyperparameters for reconstructing non-anomalous data. The focus is on the Bonn dataset, as similar findings apply to the CHB-MIT dataset. Consistent with the model selection criteria applied to both datasets, the average MSE over 10-fold cross-validation is used as the comparison metric. Only the top configurations, which showed no significant difference in MSE performance based on the hypothesis testing in the previous subsection, are considered. For reference, Table 2 provides the search space details for the Bonn dataset. The impact of each hyperparameter setting is discussed through figures, where the left panel shows the distribution of the average MSE across the 10 folds, and the right panel illustrates the frequency of each model configuration.
Latent dimension
In the selected configurations, the dimensionality of the latent space is consistently set to 32, 64, or 128, as shown in Figure 4. Lower dimensionalities, such as 8 and 16, were found to significantly degrade network performance. This observation aligns with findings in the literature on AEs [28], suggesting that excessive compression prevents the VAE’s capacity to learn fundamental non-seizure patterns needed to achieve sufficient reconstruction accuracy for the input EEG recordings. On the other hand, undercomplete AEs, where the latent dimension is smaller than the input dimension, are generally preferred as they force the encoder to extract the most salient features from the input data. However, this property diminishes as the latent dimension approaches the input dimension, leading to an overcomplete representation [28]. Since no significant difference in MSE performance was observed across the configurations with latent dimensions of 32, 64, or 128, a dimensionality of 32 is recommended for anomaly detection. This choice promotes more effective identification and representation of the critical features of non-seizure data, enhancing the model’s ability to distinguish it from anomalous seizure data.
Activation function
None of the configurations use the Tanh activation function, relying exclusively on ReLU or Identity activations, as shown in Figure 5. Notably, 112 out of 131 configurations incorporate the Identity function, consistent with the choice made by the SincNet authors in [51]. This strong preference suggests that preserving the original structure of the EEG recordings after the SincNet layer’s decomposition may enhance the VAE’s effectiveness more than applying non-linear transformations. In this scenario, the Identity function is also preferred since it introduces no additional complexity compared to non-linear transformations.
Filters
No configuration utilizes 2, 4 and 8 filters, as shown in Figure 6. This suggests that a minimal filter count is insufficient for effective feature extraction, which is crucial for the proper functioning of the VAE. Therefore, more detailed frequency band extraction is necessary to enhance performance. Among the different configurations, having 32 filters lead to slightly better performance. However, it is important to remark that, based on the hypothesis tests, no significant performance differences were observed between the configurations. Thus, the choice of the number of filters in the SincNet layer may be driven by factors such as inference speed, model complexity, or the need for a finer resolution in signal decomposition. In this work, as noted in the previous subsection, the primary goal was to maintain low model complexity, which is why 16 filters were selected.
Kernel size
None of the configurations use kernel sizes of 3, 5, 7, or 11, as shown in Figure 7. This suggests that these kernel sizes may lead to a low resolution in the Hamming windows applied by each filter during signal processing. Larger kernel sizes might offer better resolution, allowing for more detailed frequency band extraction. As for the number of filters, the choice of kernel size must ultimately be dictated by the specific needs of the task, balancing between resolution and computational efficiency.